# Steady-State Error Compensation for Reinforcement Learning with Quadratic Rewards

### v1

@@ Abstract
The selection of a reward function in Reinforcement Learning (RL) has garnered significant attention because of its impact on system performance. Issues of steady-state error often manifest when quadratic reward functions are employed. Although existing solutions using absolute-value-type reward functions partially address this problem, they tend to induce substantial fluctuations in specific system states, leading to abrupt changes. In response to this challenge, this study proposes an approach that introduces an integral term. By integrating this term into quadratic-type reward functions, the RL algorithm is adeptly tuned, augmenting the system’s consideration of long-term rewards and, consequently, alleviating concerns related to steady-state errors. Through experiments and performance evaluations on the Adaptive Cruise Control (ACC) model and lane change models, we validate that the proposed method not only effectively diminishes steady-state errors but also results in smoother variations in system states.
@@
Reinforcement Learning, Reward Function, Steady-State Error, PID.
@@ I Introduction
Reinforcement learning is a data-driven, model-free approach[1]. In 2015, Mnih et al. pioneered the fusion of deep learning with reinforcement learning by introducing the Deep Reinforcement Learning (DRL) algorithm, which significantly augmented the representational capacity and generalization performance of reinforcement learning algorithms[2]. Today, this approach has extensive applications across various industries, including gaming, robotics, natural language processing, healthcare, Industry 4.0, smart grids, and beyond[3].
@@
Optimal control theory is centered on crafting control policies to attain optimal performance for a given system based on a specific performance metric. This metric is commonly expressed as a cost or objective function, and the objective of optimal control is to discover a control policy that minimizes or maximizes this performance metric[4]. However, conventional optimization methods may encounter challenges when confronted with uncertainties and changes in dynamic environments[5]. RL serves as a bridge between traditional optimal control and adaptive control algorithms[6].
@@
In control engineering, RL is adept at uncovering optimal control laws. Through iterative interactions with the environment, optimal control strategies can be identified without necessitating a precise system model, relying instead on a trial-and-error approach. Reinforcement Learning is conceptualized as a Markov Decision Process (MDP), where the agent (controller) observes the state of the environment, takes actions, receives rewards or penalties, and refines its policy based on this feedback. Through continual learning and optimization, reinforcement learning progressively unveils the optimal control law, empowering the system to attain optimal performance in specific tasks or goals[7]. The reward function in reinforcement learning is akin to the cost function in control theory, defining the agent’s objectives in the environment and exerting direct influence over the algorithm’s performance, convergence speed, and the nature of the optimal policy[8].
As reinforcement learning algorithms advance, reconsideration of reward function design becomes imperative. Booth et al. discovered that reward functions precisely aligned with actual task performance metrics often tend to be sparse[9]. Real task metrics may encode success as a reward of 1 and failure as 0. Sparse task metrics are frequently challenging to learn, leading to the common use of dense reward functions in practice. These dense reward functions are typically crafted through trial and error by experts[10]. While dense reward functions prove more efficient in exploration, they necessitate ongoing interaction with the environment for tuning, delicately balancing the relative value of each state-action pair to ultimately derive a robust reward function[11]. The output format of the reward function, the setting of parameter weights, and the specific content of reward terms all have a significant influence on the final control performance.
@@
Engel et al. established a connection between the reward function and the ultimate control performance of closed-loop systems. Their research revealed that a reward function grounded in quadratic values could result in notable steady-state errors in the final state variables, attributed to the discounting nature of rewards in RL algorithms. Furthermore, their exploration of absolute value-based reward functions demonstrated a decrease in steady-state errors and a learning curve resembling that of quadratic rewards. However, it was noted that the learning time for absolute value-based curves significantly increased compared to quadratic rewards[12].
To guarantee the reasonable characteristics of the learned curves in reinforcement learning, we introduce the Interior-Point Optimization (IPO) method for reference. The IPO solution resides within the set of inequality constraints and closely approximates the true optimal solution. The primary contributions of this study can be summarized as follows:
@@
This study identified that the reward function based on the absolute value still induces considerable fluctuations in specific state changes throughout the entire control process, resulting in prominent peaks. Nevertheless, the introduction of an integral term offers a dual benefit, it helps diminish the steady-state error of the square function and concurrently alleviates the peaks associated with the absolute value. The subsequent sections of this paper are as follows:
@@
Section \@slowromancapii@ introduces the basic theory of reinforcement learning and the Deep Deterministic Policy Gradient (DDPG) algorithm. Section \@slowromancapiii@ outlines the specific method for introducing the integral term. Section \@slowromancapiv@ details the modeling of the experimental environment and conducts simulations to validate the proposed method. Section \@slowromancapv@ summarizes the entire paper.
@@ II-A Reinforcement Learning
A MDP serves as the ideal mathematical model for RL[13]. In a standard RL process, it is initially assumed that the environment $e$ interacts with the agent. At each time step $t$, the agent receives observations $s$ from the environment, takes action $a$, causing the environment to change on the basis of the agent’s actions, and receives an immediate reward $r$. Generally, the environment’s state is partially observable to the agent, often considering a set of historical state-action pairs as the current state. In this context, we consider reinforcement learning to be entirely observable, allowing the agent to access global information.
The agent’s actions are determined by policy $\pi$. The policy $\pi(a|s)=P(A_{t}=a|S_{t}=s)$ is a function that expresses the probability of taking action $a$ given the input state $s$. When a policy is deterministic, it outputs only one specific action with a probability of one at each state, whereas the probabilities of other actions are 0. In the case of a stochastic policy, it outputs a probability distribution over actions at each state, and action can be sampled on the basis of this distribution. Probability of the agent transitioning from state $s$ to the next state $s^{\prime}$ after taking action $a$ according to policy $\pi$ is defined as $P(s^{\prime}|s)=P(s_{t+1}=s^{\prime}|s_{t}=s)$, representing the state transition probability with Markovian properties. In a MDP, starting from time step 0 until reaching the termination state, the sum of discounted rewards is referred to as the return.
@@ II-B DDPG Algorithm
The DDPG algorithm combines concepts from deterministic policy gradient algorithms and extends the action space to continuous space, thus establishing itself as a model-free deep reinforcement learning algorithm[14]. Built upon the actor-critic framework, it employs deep neural networks to approximate both the policy network and the action-value function. Training of the policy network and value network model parameters is accomplished through stochastic gradient descent. The algorithm adopts a dual neural network architecture for both the policy and value functions, consisting of online and target networks, enhancing stability during the learning process and expediting convergence. In addition, the algorithm incorporates an experience replay mechanism, where data generated by the actor interacting with the environment is stored in an experience pool. Batch data samples are then extracted for training, similar to the experience replay mechanism in Deep Q-Network, with the aim of eliminating sample correlations and dependencies to improve convergence. The pseudocode is provided as Algorithm 1.
Algorithm 1 Pseudocode of the DDPG algorithm
   Input: Gradient stepsize $\alpha$, batch size B, discount factor
             $\gamma$, soft target update rate $\tau$, random noise $\mathcal{N}_{t}$
   Randomly initialize the critic network $Q(s,a|\omega^{Q})$ and actor
    network $\mu(s|\omega^{\mu})$ with weights $\omega^{Q}$ and $\omega^{\mu}$.
   Initialize target network $Q^{\prime}$ and $\mu^{\prime}$ with weights
    $(\omega^{Q^{\prime}}\leftarrow\omega^{Q},\omega^{\mu^{\prime}}\leftarrow\omega^{\mu})$
   Initialize replay buffer $\mathcal{D}$
   for each iteration do
        Receive initial observation state $s_{1}$
        For t=1,T do
            Select action $a_{t}=\mu(s_{t}|\omega^{\mu}+\mathcal{N}_{t})$ according to
             the current policy and exploration noise
            Execute action $a_{t}$ and observe reward $r_{t}$ and observe
             new state $s_{t+1}$
            Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in replay buffer D
            Sample a random batch size of B transitions
            $(s_{i},a_{i},r_{i},s_{i+1})$ from D
            Compute the target critic network
             $y_{i}=r_{i}+\gamma Q^{\prime}(s_{i+1},\mu^{\prime}(s_{i+1}|\omega^{\mu^{\prime}})|\omega^{Q^{\prime}})$
            Update critic by minimizing the loss: $L_{Q}$
             $\frac{1}{B}\sum_{i}(y_{i}-Q(s_{i},a_{i}|\omega^{Q}))^{2}$
            Update actor by minimizing the loss: $L_{\mu}$
             -$\frac{1}{B}\sum_{i}\nabla_{a}Q(s,a|\omega^{Q})|_{s=s_{i},a=\mu(s_{i})}\nabla_{\omega^{\mu}}\mu(s|\omega^{\mu})|_{s_{i}}$
            Update target networks:
             $\omega^{Q^{\prime}}\leftarrow\tau\omega^{Q}+(1-\tau)\omega^{Q^{\prime}}$
             ${\omega}^{\mu^{\prime}}\leftarrow{\tau\omega}^{\mu}+(1-\tau)\omega^{\mu^{\prime}}$
        end for
   end for
@@ III-A Problem Statement
In this context, $x_{t+1}$ and $u_{t}$ is the state and action, $\mathbf{x_{g}}$ represents the desired state value, and $\mathbf{L},\mathbf{M},\mathbf{N}$ are weighting matrices. These matrices are used to fine-tune the relative impacts of the state, action vector, and rate of state changes on the overall cost. Among these, the rate of state changes considers factors such as comfort and safety in specific control problems. For instance, in the context of a moving car, if the acceleration change rate $jerk$ is too high, it will make passenger uncomfortable. In contrast to the cost function in optimal control, where the objective is to minimize a cost function, reinforcement learning strives to maximize expected returns.
@@
$C_{abs}=-\left(|\mathbf{O}(\mathbf{x}_{t+1}-\mathbf{x}_{g})|+|\mathbf{P}\mathbf{u}_{t}|+|\mathbf{Q}\dot{\mathbf{x}}_{t+1}|\right)$ (6)
Where $\mathbf{O},\mathbf{P},\mathbf{Q}$ is the weight matrix. After determining the specific reward functions, experiments were conducted using both the square and absolute value reward functions, as illustrated in Section \@slowromancapiv@. The results clearly demonstrate that, for the square reward function, although it achieves relatively smooth control, the difference between $\mathbf{x}_{t+1}$ and $\mathbf{x}_{g}$ is larger than that for the absolute value reward function. On the other hand, the absolute value reward function, while achieving smaller steady-state errors, exhibits higher peak rates of fluctuation in the state during the control process.
@@ III-B Solution
The two reward functions presented in this study are both derived from the square reward function, chosen for its superior smoothness, which helps avoid significant peaks in the system’s variables. The fundamental concept is to mitigate the steady-state error associated with the square reward function. Analogous to the incorporation of integral terms in PID control to diminish steady-state errors in control systems, this study aggregates the system’s steady-state errors and integrates them as a novel penalty into the reinforcement learning reward function. The goal is to prompt the agent to prioritize the subtle yet persistent penalties embedded in the reward function.
@@
$C_{pi}^{t}=C_{qua}^{t}+C_{I}^{t}$ (7)
$C_{I}^{t}=(\mathbf{c}_{I}^{t})^{\top}\mathbf{Z}\mathbf{c}_{I}^{t}$
$\mathbf{c}_{I}$ represents the accumulated error term. Through experiments, it has been observed that if the weights of the accumulated error over the entire control process are summed up, the training process becomes very slow and sometimes struggles to converge. Therefore, two methods are proposed to calculate the accumulated error term. The first method is as follows:
$\mathbf{c}_{I}^{t+1}=\left\{\begin{array}[]{lr}\enspace\quad\quad\quad 0\quad\quad\quad\quad if\quad 0<t\leq t_{threshold}\\ \mathbf{c}_{I}^{t}+\left|\mathbf{x}_{t+1}-\mathbf{x}_{g}\right|\quad if\quad t>t_{threshold}\end{array}\right.$ (8)
Where $\left|\mathbf{x}_{t+1}-\mathbf{x}_{g}\right|$ is used to describe the error value at time $t+1$, $\mathbf{c}_{I}^{t}$ is used to describe the accumulated error from the previous time step. The variable $t_{threshold}$ is defined as a threshold, and its setting depends on when the system enters a steady state. This approach does not calculate errors before the system enters a steady state; it only sums up the smaller steady-state errors after stabilization. On the one hand, it allows for the use of relatively large weight values to reduce steady-state errors, and on the other hand, it significantly improves the convergence speed of reinforcement learning.
In most cases, if the approximate time to reach steady state is unknown in advance and cannot be obtained through experimentation, the first method may fail to determine its critical value. Hence, we introduce the second method:
$\mathbf{c}_{I}^{t+1}=\mathbf{c}_{I}^{t}+\kappa\left|\mathbf{x}_{t+1}-\mathbf{x}_{g}\right|$ (9)
Here, $\kappa$ is a weighting factor that can be artificially set depending on the environment. Its purpose is to reduce the weight of errors during transient changes, ensuring that the RL strategy is not excessively influenced by accumulated transient change errors, thus accelerating convergence. Therefore, $\kappa$ should gradually increase over time steps, aligning with our increasing emphasis on steady-state errors. For example, a direct proportionality function. In this study, we use the $\mathtt{sigmoid}$ function to represent the $\kappa$ term:
$\kappa=\frac{1}{1+ae^{T/2-t}}$ (10)
@@ IV Simulation
This section primarily introduces the two models applied in the experiments of this article: the ACC and lane change models, along with the design of their reward functions. Additionally, the presentation of the experimental results is discussed.
IV-A Model
IV-A1 ACC
The schematic of the ACC is shown as Fig. 1.
Figure 1: Schematic for two-car following[15].
Where $f$ represents the tracked target vehicle, $r$ is the controlled vehicle $b_{f}$ is the length of the target vehicle $b_{r}$ is the length of the controlled vehicle, $l_{r}$ is the distance traveled by the controlled vehicle, and $l_{f}$ is the distance traveled by the target vehicle. The equations for the spacing and velocity errors between the target and controlled vehicles are as follows:
@@ IV-A1 ACC
Where $d_{0}$ is the standstill safety distance and $h$ is the headway time. This study adopts a first-order inertia system to approximate the vehicle motion process, leading to the following state-space equations:
@@
For the RL in the ACC environment, the state variables are set as ($e,e_{v},a$), the output actions as $u_{r}$, and the initial state is configured as (5, 5, 0).
IV-A2 Lane change
The vehicle dynamics model used for lane change in this study is illustrated in Fig. 2:
Figure 2: Dynamic bicycle model for lane change[16].
The state-space equations are as follows:
$\ \dot{x}=u\cos{\varphi}-v\sin{\varphi}$ (13)
$\ \dot{y}=v\cos{\varphi}+u\sin{\varphi}$
$\ \dot{\varphi}=\omega$
$\ \dot{u}=v\omega-\frac{1}{m}F_{Y1}\sin{\delta}$
$\ \dot{v}=-u\omega+\frac{1}{m}(F_{Y1}\cos{\delta}+F_{Y2})$
$\ \dot{\omega}=\frac{1}{I_{z}}(l_{f}F_{Y1}\cos{\delta}-l_{r}F_{Y2})$
$\ \dot{a}=\frac{\delta-a}{\tau}$
Where $x$ and $y$ represent the distances of the vehicle along the $x$ and $y$ axes, $\varphi$ is the angle between the vehicle body and the x-axis, $u$ is the velocity of the vehicle center of mass along the body, $v$ is the velocity of the vehicle center of mass perpendicular to the body, $\omega$ is the angular velocity of the vehicle body, and a stores the decision $\delta$ made by the vehicle in the previous time step. $F_{Y1},F_{Y2}$ is the lateral force. under mild steering angle, we can assume that:
$\ F_{Y1}\approx k_{f}(\frac{v+l_{f}\omega}{u})-\delta$ (14)
$\ F_{Y2}\approx k_{r}\frac{v-l_{r}\omega}{u}$
The lane change scenario considered in this paper involves the simplest case, as depicted in Fig. 7. The controlled vehicle smoothly approaches the centerline of the second lane. Therefore, the vehicle’s acceleration is defined as 0, and the state variable $x$ is not considered during the control process. The distance of the vehicle in the y-direction from the centerline of the second lane is defined as $e$, which serves as a measure of the steady-state error in vehicle control.
For the RL in the lane change environment, the state variables are set as ($e,\varphi,u,v,\omega,a$), and the output actions as $\delta$. The initial state is configured as (4, 0, 30, 0, 0, 0).
IV-B Reward function

### v2

@@ Abstract
The selection of a reward function in Reinforcement Learning (RL) has garnered significant attention because of its impact on system performance. Issues of significant steady-state errors often manifest when quadratic reward functions are employed. Although absolute-value-type reward functions alleviate this problem, they tend to induce substantial fluctuations in specific system states, leading to abrupt changes. In response to this challenge, this study proposes an approach that introduces an integral term. By integrating this integral term into quadratic-type reward functions, the RL algorithm is adeptly tuned, augmenting the system’s consideration of reward history, and consequently alleviates concerns related to steady-state errors. Through experiments and performance evaluations on the Adaptive Cruise Control (ACC) and lane change models, we validate that the proposed method effectively diminishes steady-state errors and does not cause significant spikes in some system states.
@@
Reinforcement Learning, Quadratic Reward Function, Steady-State Error, PID.
@@ I Introduction
Reinforcement learning is a data-driven approach to decision-making[1]. In 2015, Mnih et al. pioneered the fusion of deep learning with reinforcement learning by introducing the Deep Reinforcement Learning (DRL) algorithm, which significantly augmented the representational capacity of reinforcement learning algorithms[2]. Today, this approach has extensive applications across various industries, including gaming, robotics, natural language processing, healthcare, Industry 4.0, smart grids, and beyond[3].
@@
Optimal control theory is centered on crafting control policies to attain optimal performance for a given system based on a specific performance metric. This metric is commonly expressed as a cost or objective function, and the objective of optimal control is to discover a control policy that minimizes or maximizes this performance metric[4].
@@
In control engineering, RL is adept at uncovering optimal control laws. Through iterative interactions with the environment, optimal control strategies can be identified, relying instead on a trial-and-error approach. Reinforcement Learning is conceptualized as a Markov Decision Process (MDP), where the agent (controller) observes the state of the environment, takes actions, receives rewards, and refines its policy based on feedback. Through learning and optimization, reinforcement learning progressively unveils the optimal control law, empowering the system to attain optimal performance in specific tasks or goals[5]. The reward function in reinforcement learning is akin to the cost function in control theory, defining the agent’s objectives in the environment and exerting direct influence over the algorithm’s performance and convergence speed[6].
@@
Engel et al. established a connection between the reward function and the ultimate control performance of closed-loop systems. Their research revealed that a reward function grounded in quadratic values could result in notable steady-state errors in the final state variables. Furthermore, their exploration of absolute value-based reward functions demonstrated a decrease in steady-state errors [7].
Weber et al. proposed an integral action state augmentation for arbitrary actor-critic based-RL controllers, which can reduce the steady-state error in reference and disturbance rejection tracking control problems. The approach doubles the number of output neurons within the actor (one additional integral term per output), and each output neuron pair will route one neuron to the environment and the other to the integrator unit. Furthermore, it uses a steady-state error compensation specific reward to enhance the learning process.[8] Different from the methods mentioned above, this paper focuses on integrating the error integral into the reward function without changing the structure of the neural network, which is easier to implement.
@@
The primary contributions of this study can be summarized as follows: This study identified that the reward function based on the absolute value still induces considerable fluctuations in specific state changes throughout the entire control process, resulting in jerky control performance. Nevertheless, the introduction of an integral term helps diminish the steady-state error of the quadratic function without inflicting significant spikes of certain system states.
@@
The subsequent sections of this paper are as follows: Section \@slowromancapii@ introduces the basic theory of reinforcement learning and the Deep Deterministic Policy Gradient (DDPG) algorithm. Section \@slowromancapiii@ outlines the specific method for introducing the integral term. Section \@slowromancapiv@ details the modeling of the experimental environment and conducts simulations to validate the proposed method. Section \@slowromancapv@ summarizes the entire paper.
@@ II-A Reinforcement Learning
An MDP serves as the ideal mathematical model for RL[9]. In a standard RL process, it is initially assumed that the environment $e$ interacts with the agent. At each time step $t$, the agent receives observations $s$ from the environment, takes action $a$, causing the environment to change on the basis of the agent’s actions, and receives an immediate reward $r$. In this context, we consider reinforcement learning to be entirely observable, allowing the agent to access global information.
The agent’s actions are determined by policy $\pi$. The policy $\pi(a|s)=P(A_{t}=a|S_{t}=s)$ is a function that expresses the probability of taking action $a$ given the input state $s$. When a policy is deterministic, it outputs only one specific action with a probability of one at each state, whereas the probabilities of other actions are 0. In the case of a stochastic policy, it outputs a probability distribution over actions at each state, and action can be sampled on the basis of this distribution. Probability of the agent transitioning from state $s$ to the next state $s^{\prime}$ after taking action $a$ according to policy $\pi$ is defined as $P(s^{\prime}|s)=P(s_{t+1}=s^{\prime}|s_{t}=s)$, representing the state transition probability with Markovian properties. In an MDP, starting from time step 0 until reaching the termination state, the sum of discounted rewards is referred to as the return.
@@ II-B DDPG Algorithm
The DDPG algorithm combines concepts from deterministic policy gradient algorithms[10]. Built upon the actor-critic framework, it employs deep neural networks to approximate both the policy network and the action-value function. Training of the policy network and value network model parameters is accomplished through batch gradient descent. The algorithm adopts a dual neural network architecture for both the policy and value functions, consisting of online and target networks, enhancing stability during the learning process and expediting convergence. In addition, the algorithm incorporates an experience replay mechanism, where data generated by the actor interacting with the environment is stored in an experience pool. Batch data samples are then extracted for training, similar to the experience replay mechanism in Deep Q-Network, with the aim of eliminating sample correlations and dependencies to improve convergence.
@@ III-A Problem Statement
In this context, $x_{t+1}$ and $u_{t}$ is the state and action, $\mathbf{x_{g}}$ represents the desired state value and is a constant conference. $\mathbf{L},\mathbf{M},\mathbf{N}$ are weight matrices. These matrices are used to fine-tune the relative impacts of the state, action vector, and rate of state changes on the overall cost. Among these, the rate of state changes considers factors such as comfort in specific control problems. For instance, in the context of a moving car, if the acceleration change rate $jerk$ is too high, it will make passenger uncomfortable. In contrast to the cost function in optimal control, where the objective is to minimize a cost function, reinforcement learning strives to maximize expected returns.
@@
$C_{abs}=-\left(\mathbf{O}|(\mathbf{x}_{t+1}-\mathbf{x}_{g})|+\mathbf{P}|\mathbf{u}_{t}|+\mathbf{Q}|\dot{\mathbf{x}}_{t+1}|\right)$ (6)
Where $\mathbf{O},\mathbf{P},\mathbf{Q}$ are the weight matrices. After determining the specific reward functions, experiments were conducted using both the quadratic and absolute value reward functions, as illustrated in Section \@slowromancapiv@ Simulation. The results clearly demonstrate that, for the quadratic reward function, although it achieves relatively smooth control, the difference between $\mathbf{x}_{t+1}$ and $\mathbf{x}_{g}$ is larger than that for the absolute value reward function at steady-state. On the other hand, the absolute value reward function, while achieving smaller steady-state errors, exhibits higher fluctuation in the state during the control process.
@@ III-B Solution
The two reward functions presented in this study are both derived from the quadratic reward function, chosen for its superior smoothness, which helps avoid significant peaks in the system’s variables. The fundamental concept is to mitigate the steady-state error associated with the quadratic reward function. Analogous to the incorporation of integral terms in PID control to diminish steady-state errors in control systems, this study aggregates the system’s steady-state errors and integrates them as a novel penalty into the reinforcement learning reward function.
@@
$C_{pi}=C_{qua}+C_{I}$ (7)
$C_{I}=(\mathbf{c}_{I})^{\top}\mathbf{Z}\mathbf{c}_{I}$
$\mathbf{c}_{I}$ represents the accumulated error term. $Z$ is the weight value. Through experiments, it has been observed that if the accumulated errors throughout the entire control process are directly summed up, the training process becomes very slow and sometimes struggles to converge. Therefore, two methods are proposed to calculate the accumulated error term.
III-B1 Method 1
The first method is as follows:
$\mathbf{c}_{I}=\left\{\begin{array}[]{lr}\enspace\quad\quad\quad\quad 0\quad\quad\quad\quad\enspace if\quad 0<t<t_{threshold}\\ \sum_{t_{threshold}}^{t}\left|\mathbf{x}_{t}-\mathbf{x}_{g}\right|\quad if\quad t\geq t_{threshold}\end{array}\right.$ (8)
Where $\left|\mathbf{x}_{t}-\mathbf{x}_{g}\right|$ is used to describe the error value at time $t$. The variable $t_{threshold}$ is defined as a threshold, and its setting depends on when the system enters a steady-state. This approach only sums up the smaller steady-state errors after stabilization. Therefore, it allows the use of relatively large weight values $Z$ to reduce steady-state errors. Meanwhile, it significantly improves the convergence speed of reinforcement learning.
III-B2 Method 2
In most cases, if the approximate time to reach steady-state is unknown in advance and cannot be obtained through experimentation, the first method may fail to determine its critical value. Hence, we introduce the second method:
$\mathbf{c}_{I}=\sum_{0}^{t}\kappa(t)\left|\mathbf{x}_{t}-\mathbf{x}_{g}\right|$ (9)
Here, $\kappa(t)$ is a weight factor that can be artificially set depending on the environment. Its purpose is to ensure that the RL strategy is not excessively influenced by accumulated transient change errors, thus accelerating convergence. Therefore, $\kappa(t)$ should gradually increase over time steps, aligning with our increasing emphasis on steady-state errors. In this study, we use the $\mathtt{sigmoid}$ function to represent the $\kappa(t)$ term[11]:
$\kappa(t)=\frac{1}{1+ae^{T/2-t}}$ (10)
@@ IV Simulation
This section primarily introduces the two models applied to validate the proposed two solutions: the ACC and lane change models. In addition, the presentation of the experimental results is discussed.
IV-A ACC
IV-A1 Model
The schematic of the ACC is shown in Fig. 1.
Figure 1: Schematic for two-car following.
Where $f$ represents the front vehicle, $r$ is the controlled vehicle, $b_{f}$ is the length of the target vehicle, $b_{r}$ is the length of the controlled vehicle, $l_{r}$ is the distance traveled by the controlled vehicle, and $l_{f}$ is the distance traveled by the target vehicle. The equations for the spacing and velocity errors between the target and controlled vehicles are as follows:
@@ IV-A1 Model
Where $d_{0}$ is the standstill safety distance and $h$ is the headway time. This study adopts a first-order inertia system to approximate the vehicle motion process, leading to the following state-space equations[12]:
@@
For the RL in the ACC environment, the state variables are set as ($e,e_{v},a_{r}$), the output actions as $u_{r}$, and the initial state is configured as (5, 5, 0).
IV-A2 Reward function